# Wireless Datasets for Aerial Networks

### v1

@@ VII-A Description of Hardware and Software
AERPAW testbed constitutes various types of wireless radio access platforms of 4G/5G. One of them is a commercial grade 5G network based on Ericsson equipment, which is the main focus of the experiment for these datasets. AERPAW infrastructure for these experiments involves a) Ericsson 5G network with RAN and Core, b) SPN, c) AERPAW fixed node (AFN) as application server of user plane traffic, and d) the UAV to carry the SPN during aerial experimentation (Fig. 10). The RAN of this 5G system is deployed at the AERPAW LWRFL, which is a rural agricultural area (see Fig. 2(a)), predominantly an open aerial field with some vegetation on the ground.
@@
For aerial experiments, we have used three types of portable nodes as 5G devices carried by UAVs based on a) 5G modem, b) Keysight Nemo device, and c) COTS Android device with PawPrints software. The portable node based on the 5G modem as shown in Fig. 11 uses a Quectel 5G module to connect to the Ericsson 5G network, a LattePanda module as a companion computer to interface with the UAV, and an ATT 5G modem for C2. The portable node that is used as UE is mounted on a small AERPAW Multi-rotor (SAM) UAV as shown in Fig. 11.
@@
Figure 11: SAM (top) mounted with a portable radio node (bottom) used in A2G measurements [41].
For Android device-based experiments, the measurement campaigns were conducted using SAM carrying a SPN with a Nemo and a PawPrints devices as the payload. Custom hardware mounts were designed for this SPN to carry two Android smartphones simultaneously during aerial flights. The two Android devices used were: a commercial standard Samsung S21 device with PawPrints, an internal AERPAW-built Android application using open source Android APIs to log radio KPIs, and a modified Samsung S23 containing Keysight Nemo software [72], with custom firmware, access to internal modem metric, and a wider range of KPIs. A Latte Panda on-board the SPN collected and logged GPS and UAV attitude measurements, obtained from GPS sensors and an Ardupilot, respectively, along with radio KPIs streamed by the Android phone over USB. These Android phones connected to the private 4G/5G Ericsson cell tower. These phones also connected as clients to the iPerf Server at the AFN, enabling throughput measurements.
@@ VII-B Dataset Format
The 5G-NSA datasets contain two different categories based on the type of portable node used: a) 5G modem-based SPN and b) Android device-based SPN.
VII-B1 5G modem datasets
This dataset provides comprehensive RF/PHY layer and throughput measurements from an Ericsson 5G-NSA network on the AERPAW platform captured during an UAV flight mission conducted on September 15 and October 26, 2023. A Portable Node equipped with a Quectel 5G Modem mounted on an UAV was used for these experiments. The UAV was programmed to follow a predefined zigzag flight path at 30 meter altitude covering the north-west quadrant of the area from the BS tower, while the portable node captures and records the logs. This data set specifically includes RF/PHY layer performance metrics such as RSRP, SINR, Channel Quality Indicator (CQI), Rank Indicator (RI), Modulation and Coding Scheme (MCS) on the LTE and NR carriers as well as the throughput, the application layer performance metric. The same experiment was repeated for different yaw positions of the UAV, 315 and 45 degrees, to understand the impact of angular orientation of portable node antennas with respect to the BS antennas. An iperf application with downlink traffic between the client at the portable node and server at the AERPAW fixed node was used for this experiment.
The data set comprises raw data collected at the portable node right from the experiment as well as a set of post-processed log files. The following four files contain the raw data logs from the sample experiment.
  • Serving_cell_Params_ENDC.csv: Serving cell configuration and RF measurements such as Cell Id, RSRP, SINR, eARFCN, band, and subcarrier spacing on both the LTE and NR carriers of the 5G NSA network
  • Basic_and_Other_Params.csv: PHY layer parameters such as downlink MCS values and the measured and reported channel state information CQI, PMI, and RI by the portable node.
  • <date>_vehicleOut.txt: The vehicle telemetry log recorded during the UAV flight that includes its geolocation, orientation, speed with date and time.
  • <date>_iperfclient_log.txt: The raw iperf3 log captured in real time during the flight contains the interval, transfer size in megabytes, and bandwidth in megabytes/sec.
The above raw log files are post-processed to generate the following CSV data files of individual RF and PHY parameters.
  • inputf1_cellid_with_header.csv: Serving cell cell-id of LTE sector of 5G-NSA.
  • inputf2_cellid_with_header.csv: Serving cell cell-id of NR sector of 5G-NSA.
  • inputf1_rsrp_with_header.csv: RSRP values of LTE carrier, recorded in dBm.
  • inputf2_rsrp_with_header.csv: RSRP values of the NR carrier, recorded in dBm. The RSRP values indicate the power level of the reference signals received from the cell.
  • inputf1_sinr_with_header.csv: SINR data of the LTE carrier, provided in dB.
  • inputf2_sinr_with_header.csv: SINR data of the NR carrier, provided in dB. The SINR values offer insights into the quality of the signal relative to background noise and interference.
  • inputf2_cqi_with_header.csv: Measured CQI values on the downlink NR carrier.
  • inputf2_mcs_with_header.csv: Downlink MCS values on the NR carrier.
  • inputf2_ri_with_header.csv: Measured RI values on the downlink NR carrier.
  • input_throughput_with_header.csv: Downlink throughput, expressed in Mbps, the quantify the data transmission rate achieved during the UAV’s flight.
The above CSV log data include the recorded date, time and the UAV geolocation (longitude, latitude, and altitude) data. These datasets from these experiments are publicly available at the AERPAW datasets webpage [9] as well as in Dryad repository, an open data publishing platform.
((a))
((b))
((c))
((d))
Figure 12: (a) RSRP and (c) SINR of LTE and NR carriers vs. distance and time, (b) RSRP and (d) SINR of NR Carrier with UAV geo-location [41]
VII-B2 Nemo and PawPrints datasets
Depending on the setup, the Nemo and PawPrints datasets can contain up to three logs per measurement campaign: 4G/LTE logs, 5G/NR logs, and iPerf throughput logs. Within each log type, radio or throughput KPIs are merged with UAV location based on timestamps into a single CSV file. The PawPrints 4G KPIs, obtained from Android APIs, include LTE RSRP, RSRQ, Received Signal Strength Indication (RSSI), Physical Cell Identity (PCI), tracking area code (TAC), and cell id (CI). The PawPrints 5G KPIs consist of the 5G synchronization signal’s RSRP, RSRQ and RSSI. Throughput measurements are also shared whenever applicable. Thus, the PawPrints dataset consists of three files:
  • pawprints_4G_LTE.csv: PawPrints 4G log
  • pawprints_5G_NR.csv: PawPrints 5G log
  • pawprints_iperf_throughput.csv: iPerf client throughput.
The Nemo KPIs are more comprehensive and additionally include 5G MCS, CQI, and channel rank. The Nemo dataset consists of:
  • nemo_5G_NR.csv: Nemo 5G log
  • nemo_4G_LTE.csv: Nemo 4G log
@@ VII-C Representative Results
VII-C1 5G modem dataset results
As explained above, we used an UAV powered by a 5G Quectel modem to measure and collect data along a zigzag aerial path trajectory. From these experiments the RF/PHY parameters as well as application layer performance parameters were collected for analysis.
The RSRP and SINR are some of the RF parameters and CQI, RI and MCS are some of the PHY layer parameters measured and logged on the 5G-powered portable node presented here. Fig. 12(a) and Fig. 12(b) show the measured levels of downlink RSRP and SINR on LTE and NR carriers versus distance and time, whereas Fig. 12(c) and Fig. 12(d) show the same RSRP and SINR of NR carrier along the geo locations of the zigzag UAV flight trajectory. Similarly, Fig. 13 shows the reported CQI, RI and MCS versus the distance, time and geo location on the LTE and NR downlink carriers. An iperf client-server app with downlink traffic was used between the portable node and the wired AFN, and Fig. 14 shows the downlink throughput achieved at the application layer in real-time.
((a))
((b))
((c))
((d))
Figure 13: (a) CQI and (c) MCS of NR carrier with respect to distance and time, (b) CQI and (d) MCS of NR Carrier with UAV geo-location
((a))
((b))
Figure 14: (a) Throughput with respect to distance and time, (b) Throughput with UAV geo-location [41].
((a))
((b))
((c))
Figure 15: Wireless KPI data gathered at AERPAW Lake Wheeler tower 1: (a) and (b) show iPerf throughput observed by Nemo and PawPrints, respectively, while simultaneously sending traffic as iPerf clients; (c) shows the 5G downlink channel rank as recorded by Nemo.
VII-C2 Nemo and PawPrints dataset results
This section presents the representative results from the wireless KPI datasets collected from a UAV in two scenarios: measurements of a private BS at the AERPAW LWRFL with controlled UAV trajectories (using PawPrints and Nemo), and measurements of commercial cell towers from a tethered Helikite during the Packapalooza 2023 festival (using PawPrints). Fig. 15 depicts some representative results from the first scenario, when the UAV traces sawtooth trajectories in the horizontal plane, at increasing distances from the private Ericsson BS. Fig. 15(a) and Fig. 15(b) show a heatmap of the iPerf throughput measured by the Nemo and PawPrints device, respectively, when both were operating as client simultaneously. The heatmaps confirm the reduction in throughput observed near handover regions, particularly at sector boundaries, as evidenced by the blue-shaded areas. Fig. 15(c) shows the channel rank of the physical downlink channel recorded by Nemo during the flight.
@@ VII-C2 Nemo and PawPrints dataset results
The data from these experiments are publicly available at the AERPAW datasets webpage [9] and in Dryad research repository [37, 38, 39, 40].
Figure 16: RSRP variation against altitude observed by a PawPrints device on a Helikite at Packapalooza 2025.
@@ VII-D Possible Uses of Dataset
VII-D1 Study of wireless channel propagation in rural aerial environments
The empirical data on variations in signal strength, with distance from the BS and elevation and azimuth angles, can be used to validate existing theoretical propagation models and ray-tracing simulators, along with developing new mathematical or machine learning models. Moreover, deeper insights can be obtained by studying the variation in channel rank and channel quality index in aerial scenarios.
VII-D2 Application layer throughput prediction in rural aerial scenarios
Throughput prediction models for rural aerial scenarios can be created by studying the relation between physical layer KPIs such as channel rank and received SINR, and iPerf throughput.
VII-D3 Analysis of commercial cellular network coverage in aerial urban environments
The Packapalooza dataset contains received signal strength and quality values at various altitudes, which can be used to study the suitability of commercial cellular networks, designed to serve ground users, for aerial operations.
@@ VIII LoRa Propagation Dataset
Figure 17: LoRaWAN gateway and a USB-compliant programmable LoRa device, used in the AERPAW system for mobile LoRa experiments.
Figure 18: Ground and aerial vehicles that can carry LoRa devices for mobile experiments.
@@ VIII-A Description of Hardware and Software
The LoRa infrastructure consists of USB-compliant programmable LoRa devices and seven LoRaWAN gateways (see Fig. 17), which relay LoRa communications over the AERPAW backplane to tenant-dedicated Docker containers for executing application-specific data processing tasks. Standard software stacks are installed on the containers, including Prometheus and PostGRES storage, which feed data for Grafana-based visualization. These LoRa devices can be driven by ground vehicles, or carried as a payload by UAVs or a Helikite, shown in Fig. 18, thus supporting ground, aerial, and static experiments. The LoRa devices allow the experiments to configure the spreading factor between 7 to 12, and accordingly set the transmission data rate from DR3 (5.47 kbps) to DR0 (0.25 kbps).
@@ VIII-B Dataset Format
The dataset includes logs from the LoRa transmitter device and the LoRaWAN gateways. The transmitter logs include the following measurements:
  • Packet metadata such as packet ID and packet sequence number.
  • LoRaWAN datarate
  • Transmission bandwidth.
  • Code rate of transmission.
  • Timestamp of packet transmission.
  • Frequency of transmission
  • Geographic location (latitude, longitude, altitude) and orientation (yaw, pitch, roll) of the vehicle carrying the LoRa transmitter.
  • Velocity of the vehicle carrying the LoRa transmitter.
  • Transmission spreading factor.
  • GPS metadata, such as the number of satellites locked to.
The logs shared from the LoraWAN gateway include:
  • Received signal strength of the packet, as RSSI.
  • Received signal to noise ratio of the packet.
  • Timestamp of packet reception.
  • Frequency channel in which the packet was received.
  • ID of the RF chain used to receive the packet.
  • Geographic location and ID of the receiving LoRaWAN gateway.
Additionally, logs of failed transmissions are included in a CSV file, named failed_tx_packages.csv. The summary of data rates, grouped by the gateway, are also included in a CSV file, named gateway-dataRate-Table.csv. Raw signal data, containing signal-to-noise ratio (SNR) and RSSI, is shared in sigMF format. Examples of Python scripts are provided to convert from sigMF format to CSV format and to generate plots from the resultant CSV file.
@@ VIII-C Representative Results
Figure 19: Location of the LoRa tethered Helikite flight, showing location of the six LoRaWAN gateways and the top view of the Helikite locations.
The LoRa transmitter device was carried as a payload by a tethered HeliKite, which traced a free-floating trajectory over the North Campus of NC State University, Raleigh. The LoRa device was controlled by a Latte Panda mini computer, which also logged packet transmission details. The LoRa device transmitted packets, containing a sequence number and timestamp, at intervals of $1.5$ s to six LoRaWAN gateways at multiple locations, as shown in Fig. 19. These gateways recorded the received RSSI, received SNR, and the timestamp of reception. The CDF of received RSSI at the gateways is shown in Fig. 20(a). The variation in received SNR with received RSSI, over all the LoRaWAN gateways, is shown in Fig. 20(b), which indicates that higher RSSI results in a smaller range of SINR values. In contrast, lower RSSI values can cause SIR fluctuations as noise and interference dominate. The dataset not only provides physical layer signal strength parameters but also includes packet-level latency metrics, the distribution of which is shown for each transmitted packet in Fig. 20(c), aggregated over all LoRaWAN gateways.
@@
((a))
((b))
((c))
Figure 20: The KPIs of LoRaWAN links measured during a tethered Helikite experiment, where the LoRa transmitter was elevated, and data were collected across six gateways. (a) Empirical CDF of received RSSI values observed at multiple gateways. (b) Relationship between received RSSI and corresponding SNR, highlighting link quality variation. (c) End-to-end packet latency measurements, from transmission to reception, aggregated across all gateways.

### v2

@@ VII-A Description of Hardware and Software
AERPAW provides multiple wireless radio access platforms for 4G and 5G experimentation, including a commercial-grade Ericsson 5G network that serves as the primary infrastructure for these datasets. AERPAW infrastructure for these experiments involves a) Ericsson 5G network with RAN and Core, b) SPN, c) AERPAW fixed node (AFN) as application server of user plane traffic, and d) the UAV to carry the SPN during aerial experimentation (Fig. 8). The RAN of this 5G system is deployed at the AERPAW LWRFL, which is a rural agricultural area (see Fig. 2b), predominantly an open aerial field with some vegetation on the ground.
@@
(a) SAM mounted with portable node (SPN)
(b) Small portable node (SPN) with Quectel 5G modem
Figure 9: SAM mounted with a portable radio node used in A2G measurements. (a) SAM mounted with SPN and (b) SPN with Quectel 5G modem.
Figure 10: Software functional workflow with 5G-NSA network. The workflow highlights that KPIs are derived from time-synchronized radio, traffic, and UAV telemetry, which is essential for statistically reliable spatial performance mapping.
For aerial experiments, three types of portable nodes were used as 5G user equipment carried by UAVs: a modem-based node, a Keysight Nemo device, and a COTS Android device running PawPrints software. The portable node based on the 5G modem as shown in Fig. 9 uses a Quectel 5G module to connect to the Ericsson 5G network, a LattePanda module as a companion computer to interface with the UAV, and an ATT 5G modem for C2. The portable node that is used as UE is mounted on a small AERPAW Multi-rotor (SAM) UAV.
@@
(a) RSRP of LTE and NR signals versus distance and time.
(b) SINR of LTE and NR signals versus distance and time.
(c) RSRP of NR signals over 3D trajectory.
(d) SINR of NR carrier
Figure 11: (a) RSRP and (b) SINR of LTE and NR carriers vs. distance and time, (c) RSRP and (d) SINR of NR Carrier with UAV geo-location. These figures show that both RSRP and SINR exhibit clear distance- and location-dependent trends, supporting spatially resolved statistical coverage analysis.
For Android-based measurements, the SAM UAV carried an SPN payload containing two Android smartphones mounted on custom hardware brackets. The two Android devices used were: a commercial standard Samsung S21 device with PawPrints, an internal AERPAW-built Android application using open source Android APIs to log radio KPIs, and a modified Samsung S23 containing Keysight Nemo software [42], with custom firmware, access to internal modem metric, and a wider range of KPIs. A Latte Panda on-board the SPN collected and logged GPS and UAV attitude measurements, obtained from GPS sensors and an Ardupilot, respectively, along with radio KPIs streamed by the Android phone over USB. These Android phones connected to the private 4G/5G Ericsson cell tower. These phones also connected as clients to the Internet Protocol Performance (iPerf) Server at the AFN, enabling throughput measurements.
@@ VII-B Dataset Format
The 5G-NSA datasets are categorized based on the type of portable node used during the measurement campaigns: (a) 5G modem-based SPN and (b) Android device-based SPN (Nemo and PawPrints). The datasets provide synchronized RF/PHY-layer performance metrics, application-layer throughput measurements, and UAV telemetry collected during controlled UAV flight experiments on the AERPAW platform.
The modem-based datasets include LTE and NR KPIs such as RSRP, SINR, Channel Quality Indicator (CQI), Rank Indicator (RI), Modulation and Coding Scheme (MCS), and downlink throughput obtained using a UAV-mounted Quectel 5G modem connected to an Ericsson 5G-NSA network. Complementary Android-based datasets collected using Nemo and PawPrints tools provide LTE and NR KPIs derived from Android APIs, along with throughput measurements when available.
All datasets are released as structured CSV files with synchronized timestamps and UAV geolocation (longitude, latitude, altitude), enabling spatially resolved analysis of aerial cellular performance. Detailed descriptions of file organization, raw and post-processed logs, and parameter-specific data files for both modem-based and Android-based datasets are provided in Appendix A.
@@ VII-C Representative Results
5G modem dataset results: As explained above, we used an UAV powered by a 5G Quectel modem to measure and collect data along a zigzag aerial path trajectory. From these experiments the RF/PHY parameters as well as application layer performance parameters were collected for analysis.
The RSRP and SINR are some of the RF parameters and CQI, RI and MCS are some of the PHY layer parameters measured and logged on the 5G-powered portable node presented here. Fig. 11a and Fig. 11b show the measured levels of downlink RSRP and SINR on LTE and NR carriers versus distance and time, whereas Fig. 11c and Fig. 11d show the same RSRP and SINR of NR carrier along the geo locations of the zigzag UAV flight trajectory. Similarly, Fig. 12 shows the reported CQI, RI and MCS versus the distance, time and geo location on the LTE and NR downlink carriers. An iPerf client-server app with downlink traffic was used between the portable node and the wired AFN, and Fig. 13 shows the downlink throughput achieved at the application layer in real-time.
(a) CQI & RI of NR carrier
(b) CQI of NR carrier
(c) MCS & RI of NR carrier
(d) MCS of NR carrier
Figure 12: (a) CQI and (c) MCS of NR carrier with respect to distance and time, (b) CQI and (d) MCS of NR Carrier with UAV geo-location. The adaptation metrics vary consistently with distance and position, indicating that link adaptation responds predictably to aerial channel conditions.
(a) Throughput vs distance & time
(b) Throughput vs geo location
Figure 13: (a) Throughput with respect to distance and time, (b) Throughput with UAV geo-location. The throughput plots show significant spatial and temporal variability, highlighting that achievable rate statistics are strongly trajectory dependent.
(a) Nemo iPerf throughput.
(b) PawPrints iPerf throughput.
(c) 5G downlink channel rank (Nemo).
Figure 14: Wireless KPI data gathered at AERPAW Lake Wheeler tower 1: (a) and (b) show iPerf throughput observed by Nemo and PawPrints, respectively, while simultaneously sending traffic as iPerf clients; (c) shows the 5G downlink channel rank as recorded by Nemo. The comparison demonstrates that measured throughput statistics depend on the logging tool, which is practically important when benchmarking or cross-validating datasets.
Nemo and PawPrints dataset results: This section presents the representative results from the wireless KPI datasets collected from a UAV in two scenarios: measurements of a private BS at the AERPAW LWRFL with controlled UAV trajectories (using PawPrints and Nemo), and measurements of commercial cell towers from a tethered Helikite during the Packapalooza 2023 festival (using PawPrints). Fig. 14 depicts some representative results from the first scenario, when the UAV traces sawtooth trajectories in the horizontal plane, at increasing distances from the private Ericsson BS. Fig. 14a and Fig. 14b show a heatmap of the iPerf throughput measured by the Nemo and PawPrints device, respectively, when both were operating as client simultaneously. The heatmaps confirm the reduction in throughput observed near handover regions, particularly at sector boundaries, as evidenced by the blue-shaded areas. Fig. 14c shows the channel rank of the physical downlink channel recorded by Nemo during the flight.
@@
The data from these experiments are publicly available at the AERPAW datasets webpage [3] and in Dryad research repository [10, 83, 82, 80, 81].
Figure 15: RSRP variation against altitude observed by a PawPrints device on a Helikite at Packapalooza 2023.
@@ VII-D Possible Uses of Dataset
Study of wireless channel propagation in rural aerial environments: The empirical data on variations in signal strength, with distance from the BS and elevation and azimuth angles, can be used to validate existing theoretical propagation models and ray-tracing simulators, along with developing new mathematical or machine learning models. Moreover, deeper insights can be obtained by studying the variation in channel rank and channel quality index in aerial scenarios.
Application layer throughput prediction in rural aerial scenarios: Throughput prediction models for rural aerial scenarios can be created by studying the relation between physical layer KPIs such as channel rank and received SINR, and iPerf throughput.
Analysis of commercial cellular network coverage in aerial urban environments: The Packapalooza dataset contains received signal strength and quality values at various altitudes, which can be used to study the suitability of commercial cellular networks, traditionally designed to serve ground users, for aerial operations.
@@ VIII-A Description of Hardware and Software
The LoRa infrastructure consists of USB-compliant programmable LoRa devices and seven LoRaWAN gateways (see Fig. 16), which relay LoRa communications over the AERPAW backplane to tenant-dedicated Docker containers for executing application-specific data processing tasks. Standard software stacks are installed on the containers, including Prometheus and PostGRES storage, which feed data for Grafana-based visualization. Fig. 17 shows the mobile platforms used for LoRaWAN measurements, including a ground vehicle (Fig. 17a), a UAV (Fig. 17b), and a helikite (Fig. 17c). The LoRa devices allow the experiments to configure the spreading factor between 7 to 12, and accordingly set the transmission data rate from DR3 (5.47 kbps) to DR0 (0.25 kbps).
@@ VIII-B Dataset Format
The LoRa propagation dataset includes measurement logs collected from both the LoRa transmitter and the associated LoRaWAN gateways, enabling end-to-end characterization of aerial and ground LoRaWAN links. The dataset captures detailed packet-level metadata together with synchronized vehicle mobility and positioning information, supporting coverage analysis, signal quality evaluation, and latency assessment in aerial IoT scenarios.
Transmitter-side logs record packet identifiers, transmission parameters (including data rate, bandwidth, code rate, spreading factor, and carrier frequency), transmission timestamps, and vehicle state information such as geographic location, orientation, velocity, and GPS status. Gateway-side logs capture reception metrics for successfully decoded packets, including Received signal strength indicator (RSSI), signal-to-noise ratio (SNR), reception timestamps, frequency channels, RF chain identifiers, and gateway location metadata.
In addition to successful transmissions, the dataset includes records of failed packet transmissions and aggregated summaries of gateway-level data rate statistics, enabling reliability and performance analysis across deployment scenarios. Raw signal-level measurements are provided in SigMF format, while processed packet- and gateway-level records are released in structured CSV files suitable for statistical analysis and machine learning applications. Detailed file inventories, directory organization, and conversion script descriptions are provided in Appendix A.
(a) LoRaWAN gateway.
(b) LoRaWAN dongle.
Figure 16: (a) LoRaWAN gateway and (b) USB-compliant programmable LoRa device, used in the AERPAW system for mobile LoRa experiments.
(a) Ground vehicle carrying a LoRaWAN device.
(b) UAV carrying a LoRaWAN device.
(c) Helikite carrying a LoRaWAN device.
Figure 17: Ground and aerial platforms used in mobile LoRaWAN experiments: (a) ground vehicle, (b) UAV, and (c) helikite, each carrying a LoRaWAN device.
@@ VIII-C Representative Results
Figure 18: Location of the LoRa tethered Helikite flight, showing location of the six LoRaWAN gateways and the top view of the Helikite locations. The spatial separation of gateways enables statistical evaluation of gateway diversity and coverage overlap.
The LoRa transmitter device was carried as a payload by a tethered HeliKite, which traced a free-floating trajectory over the North Campus of NC State University, Raleigh. The LoRa device was controlled by a Latte Panda mini computer, which also logged packet transmission details. The LoRa device transmitted packets, containing a sequence number and timestamp, at intervals of $1.5$ s to six LoRaWAN gateways at multiple locations, as shown in Fig. 18. These gateways recorded the received RSSI, received SNR, and the timestamp of reception. The CDF of received RSSI at the gateways is shown in Fig. 19a. The variation in received SNR with received RSSI, over all the LoRaWAN gateways, is shown in Fig. 19b, which indicates that higher RSSI results in a smaller range of SINR values. In contrast, lower RSSI values can cause SIR fluctuations as noise and interference dominate. The dataset not only provides physical layer signal strength parameters but also includes packet-level latency metrics.
@@
(a) CDF of received RSSI at various LoRaWAN gateways.
(b) Variation in LoRa packet’s received SNR as a function of received RSSI, aggregated over six LoRaWAN gateways.
Figure 19: The KPIs of LoRaWAN links measured during a tethered Helikite experiment, where the LoRa transmitter was elevated, and data were collected across six gateways. (a) Empirical CDF of received RSSI values observed at multiple gateways. (b) Relationship between received RSSI and corresponding SNR, highlighting link quality variation.